# Yangming Zhang*    Wenqi Jia* Dept. of Computer Science

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a mainstream for novel view synthesis, leveraging continuous aggregations of Gaussian functions to model scene geometry. However, 3DGS suffers from substantial memory requirements to store the multitude of Gaussians, hindering its practicality. To address this challenge, we introduce GaussianSpa, an optimization-based simplification framework for compact and high-quality 3DGS. Specifically, we formulate the simplification as an optimization problem associated with the 3DGS training. Correspondingly, we propose an efficient “optimizing-sparsifying” solution that alternately solves two independent sub-problems, gradually imposing strong sparsity onto the Gaussians in the training process. Our comprehensive evaluations on various datasets show the superiority of GaussianSpa over existing state-of-the-art approaches. Notably, GaussianSpa achieves an average PSNR improvement of 0.9 dB on the real-world Deep Blending dataset with 10$\times$ fewer Gaussians compared to the vanilla 3DGS. Our project page is available at https://noodle-lab.github.io/gaussianspa.
Figure 1: We present GaussianSpa, enabling high-quality and compact view synthesis with superior rendering of details. Compared to the existing state-of-the-art method, Mini-Splatting [20], our GaussianSpa captures detail-rich textures and background more accurately with fewer Gaussians.
@@ 1 Introduction
Novel view synthesis has become a pivotal area in computer vision and graphics, driving advancements in applications such as virtual reality, augmented reality, and immersive media experiences [21]. NeRF [35] has recently gained prominence in this domain because it can generate high-quality, photorealistic images from sparse input views by representing scenes as continuous volumetric functions based on neural networks. However, NeRF often requires extensive computational resources and long training times, making it less practical for real-time applications and large-scale reconstructions.
@@
3D Gaussian Splatting (3DGS) [25] has emerged as a powerful alternative, leveraging continuous aggregations of Gaussian functions to model scene geometry and appearance. Unlike NeRF, which relies on neural networks to approximate volumetric radiance fields, 3DGS directly represents scenes using a collection of Gaussians. This approach excels in capturing details and smooth transitions, offering significant advantages in training and rendering speed. 3DGS achieves superior visual fidelity [28] compared to NeRF while reducing computational overhead, making it more suitable for interactive applications that demand quality and performance.
@@
Despite its strengths, 3DGS suffers from significant memory requirements that hinder its practicality. The main issue is the massive memory consumption associated with storing a large number of Gaussians that are needed to represent complex scenes. Each Gaussian occupies memory space for its parameters, including position, covariance, and color attributes. In densely sampled scenes, the sheer volume of Gaussians leads to memory usage that exceeds the capacity of typical hardware, making it challenging to handle higher-resolution scenes and limiting its applicability in resource-constrained environments.
@@
Existing works, e.g., Mini-Splatting [20], LightGaussian [19], LP-3DGS [54], EfficientGS [31], and RadSplat [40], have predominantly focused on mitigating this issue by removing a certain number of Gaussians. Techniques such as pruning and sampling aim to discard unimportant Gaussians based on hand-crafted criteria such as opacity [25, 53], importance score (hit count) [20, 19, 40], dominant primitives [31], and binary mask [54, 29]. However, these criteria are generally utilized to determine the importance of Gaussian points from a single heuristic perspective, limiting robustness in dynamic scenes or under varying lighting conditions. Moreover, the sudden one-shot removal may cause permanent loss of Gaussians that are crucial to visual synthesis, making it challenging to recover the original performance after even long-term training, as shown in Figure 2. Consequently, while these methods can alleviate memory and storage burdens to some extent, they often lead to sub-optimal rendering outcomes with loss of details and visual artifacts, thereby compromising the quality of the synthesized views.
@@
In this paper, we present an optimization-based simplification framework, GaussianSpa, for compact and high-quality Gaussian Splatting. In the proposed framework, we formulate 3DGS simplification as a constrained optimization problem under a target number of Gaussians. Then, we propose an efficient “optimizing-sparsifying" solution for the formulated problem by splitting it into two simple sub-problems that are alternately solved in the “optimizing" step and the “sparsifying" step. Instead of permanently removing a certain number of Gaussians, GaussianSpa incorporates the “optimizing-sparsifying” algorithm into the training process, gradually imposing a substantial sparse property onto the trained Gaussians. Hence, our GaussianSpa can simultaneously enjoy maximum information preservation from the original Gaussians and a desired number of reduced Gaussians, providing compact 3DGS models with high-quality rendering. Overall, our contributions can be summarized as follows:
@@
  • We comprehensively evaluate GaussianSpa through extensive experiments on various complex scenes, demonstrating improved rendering quality compared to existing approaches. Particularly, with as high as 10$\times$ fewer number of Gaussians than the vanilla 3DGS, GaussianSpa achieves an average 0.4 dB improvement on the Mip-NeRF 360 [4] and Tanks&Temples [27] datasets, 0.9 dB on Deep Blending [23] dataset. Furthermore, we conduct various visual quality assessments, showing that GaussianSpa exhibits high-quality rendering of details and sparse 3D Gaussian views.
@@ 2.1 Novel View Synthesis
Novel View Synthesis (NVS) [2] is a technique in computer vision and graphics to generate images of a 3D scene from unseen viewpoints based on existing multi-view data. A seminal work in this field is NeRF [35], which uses an implicit neural network to represent a scene as a continuous 5-dimensional function, mapping 3D coordinates and viewing directions to color and density. NeRF introduces positional encoding and a hierarchical sampling strategy to improve the ability of the MLP [6] to represent high-frequency image details and accelerate the rendering process. However, NeRF’s reliance on single narrow-ray sampling poses challenges in addressing aliasing and blur artifacts. To mitigate this, Mip-NeRF [3] represents the scene at a continuous-valued scale, enabling anti-aliased rendering by casting conical frustums from each pixel at various scales. Additionally, Mip360 [4] extends NeRF’s capabilities to unbounded scenes, utilizing a non-linear scene parameterization, online distillation, and a novel distortion-based regularizer. More recently, an explicit point-based method, 3D Gaussian Splatting (3DGS), which leverages continuous aggregations of Gaussian functions to model scene geometry, has emerged as a point-based alternative, demonstrating significant improvements in rendering quality and speed.
Figure 2: PSNR curves of human-crafted criteria-based pruning methods. All methods remove 85% of the Gaussians at iteration 25K.
@@ 2.2 Efficient 3DGS
3DGS significantly improves rendering quality and computational efficiency compared to traditional methods such as NeRF [35], which are often resource-intensive and computationally slow. However, the demands of real-time rendering and the reconstruction of unbounded, large-scale 3D scenes underscore the requirement for model efficiency [10]. Optimization for 3DGS focuses on four main categories to effectively reduce model size: Gaussian simplification, vector quantization, spherical harmonics optimization, and hybrid compression.
@@
Gaussian Simplification. Gaussian simplification involves removing Gaussians to compress 3DGS models into compact forms, significantly reducing memory costs. Prior approaches [20, 40, 34, 54, 32, 46, 30, 13, 26] mainly focus on pruning redundant Gaussians according to hand-crafted importance criteria. Mini-Splatting [20] introduces blur splitting and depth reinitialization to address overlapping and under-reconstructed problems, using stochastic sampling to better preserve geometry based on importance scores. Radsplatting [40] computes an importance score by aggregating ray contributions of Gaussians across all input images, replacing summation with a max operator for robustness to varying scene coverage. Taming 3DGS [34] implements steerable densification by calculating scores based on pixel saliency, gradients, and Gaussian attributes, then selectively cloning or splitting high-scoring Gaussians. LP-3DGS [54] applies a trainable binary mask with Gumbel-Sigmoid to automatically optimize pruning, improving efficiency while maintaining quality.
@@
Vector Quantization. Vector quantization [50] aims to compress data using discrete entries from a learned codebook, reducing data redundancy. Several works [22, 38, 12, 51, 42, 36, 39] apply vector quantization for compressing 3DGS models. For instance, EAGLES [22] quantizes attributes like spherical harmonics (SH) coefficients and opacity to create latent representations for per-point compression, using an MLP decoder to reconstruct the original attributes with high fidelity. CompGS [38] applies a K-means-based vector quantization approach to compress 3DGS parameters, significantly reducing storage and rendering time with minimal quality loss. RDO-Gaussian [51] performs end-to-end rate-distortion optimization for 3D Gaussian Splatting, combining dynamic pruning and entropy-constrained vector quantization to reduce storage while maintaining quality.
@@
Spherical Harmonics Optimization. Among 3DGS parameters, spherical harmonics (SH) coefficients take up a significant portion, requiring (45+3) floating-point values per Gaussian for a degree of 3 (as used in [25]), which accounts for around 80% of the total attribute volume. Prior works [19, 29] optimize SH coefficients to reduce storage overhead. For instance, EfficientGS [31] retains zeroth-order SH for all Gaussians, increasing the order when necessary.
@@
Hybrid Compression. In addition to methods focused on individual aspects of 3DGS compression, several recent approaches [19, 31, 29, 52] combine multiple techniques for hybrid compression to achieve further storage reduction. LightGaussian [19] prunes 3DGS points using a global importance score, applies vector quantization to compress SH coefficients, and reduces SH degree through knowledge distillation combined with pseudo-view augmentation. EfficientGS [31] not only keeps zeroth-order SH for all Gaussians, increasing the order only when needed but also classifies Gaussians into steady and non-steady states, densifying non-steady Gaussians while retaining only the top-K most important Gaussians per ray. CompactGaussian [29] introduces a learnable mask to mark redundant Gaussians and residual vector quantization for scale and rotation and replaces SH with a grid-based neural field to represent colors.
@@
Although these methods improve 3DGS efficiency, they lead to irreversible information loss during the Gaussian simplification process. Specifically, those methods perform simplification by a one-shot pruning on the Gaussians without prominent sparse characteristics, causing a notable PSNR degradation that is challenging to recover, as shown in Figure 2. Consequently, when applying additional compressions like SH optimization and vector quantization after simplification, the compressed 3DGS model often exhibits further performance loss due to the accumulated errors. Our work focuses on Gaussian simplification by formulating it as an optimization problem and imposing substantial sparsity onto Gaussians, mitigating performance issues with improved rendering quality at the first stage to achieve efficient 3DGS.
Figure 3: Overall workflow of our proposed GaussianSpa framework.
@@ 3.1 Background of 3D Gaussian Splatting
where $\bm{x}$ is an arbitrary location in the 3D scene. The covariance matrix $\bm{\Sigma}$ is generally composed of $\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{\top}\bm{R}^{\top}$ with a rotation matrix $\bm{R}$ and a scale matrix $\bm{S}$, ensuring the positive definite property. Additionally, each Gaussian is associated with an opacity $a$ and spherical harmonics (SH) coefficients.
@@
In the training process [25], the Gaussians are initialized from a sparse point cloud generated by Structure-from-Motion (SfM) techniques [47]. Then, the number of Gaussians is adjusted by a densification algorithm [25] according to the reconstruction geometry. Afterward, each Gaussian’s variables, including the center position, rotation matrix, scaling matrix, opacity, and SH coefficients, are optimized by minimizing the reconstruction errors using the standard gradient descent. Specifically, the loss $\mathcal{L}$ is given by
$\mathcal{L}=(1-\rho)\mathcal{L}_{1}+\rho\mathcal{L}_{\text{D-SSIM}},$ (4)
@@ 3.2 Problem Formulation: Simplification as Constrained Optimization
However, the above problem cannot be solved because the Gaussians are continuous functions instead of differentiable variables. Recalling the rendering process with Eq. 3, the contribution of the $i$-th Gaussians to a particular pixel is finally determined by its opacity $a_{i}$. Alternatively, the original sparsity constraint directly on the Gaussians can be explicitly transformed into a sparsity constraint on the opacity variables. As each Gaussian only has one opacity variable, we can use a vector $\bm{a}$ to represent all the opacities and easily restrict its number of non-zeros with $\ell_{0}$ norm, $\|\bm{a}\|_{0}$. For better presentation, we use a variable set $\bm{\Theta}$ to represent other GS variables except for the opacity $\bm{a}$. Hence, the constrained optimization problem can be reformulated as
@@
$h(\bm{a})=\left\{\begin{matrix}0,&\|\bm{a}\|_{0}\leq\kappa,\\ +\infty,&\textup{otherwise},\\ \end{matrix}\right.$ (7)
@@
$\min_{\bm{a},\bm{\Theta}}~{}\mathcal{L}(\bm{a},\bm{\Theta})+h(\bm{a}).$ (8)
Here, the first term is the original 3DGS training objective, which is easy to handle, while the second term is still non-differentiable. This makes the above problem unable to solve using gradient descent under existing training frameworks such as PyTorch [44] and TensorFlow [1]. To decouple the two terms, we further introduce an auxiliary variable $\bm{z}$, whose size is identical to $\bm{a}$, and convert Eq. 8 as
@@
respectively. Here, $\eta$ is the learning rate defined in the training process.
Remark-1. The “optimizing" step essentially optimizes over the original Gaussians and pushes the opacity $\bm{a}$ close to the auxiliary $\bm{z}$ that is exactly sparse using a gradient-based approach. This way, it can simultaneously satisfy the 3DGS performance and the sparsity requirements, achieving a sweet point in the training process.
[…]
$\bm{z}\leftarrow\mathrm{\bf{prox}}_{h}(\bm{a}+\bm{\lambda}).$ (16)
The proximal operator maps the opacity $\bm{a}$ to the auxiliary variable $\bm{z}$, which contains at most $\kappa$ non-zeroes. This operator results in a projection solution [8], sparsifying $(\bm{a}+\bm{\lambda})$, by projecting it onto a set that satisfies $\|\bm{z}\|_{0}\leq\kappa$. More specifically, the projection keeps top-$\kappa$ elements and sets the rest to zeros.
Remark-2. The “sparsifying" step essentially prunes the auxiliary variable $\bm{z}$, which can be considered the exactly sparse version of $\bm{a}$. By not directly operating the original Gaussians, this step avoids the irreversible loss of important information.
[…]
$\bm{\lambda}\leftarrow\bm{\lambda}+\bm{a}-\bm{z}.$ (17)
Overall, the optimization steps and the multiplier update are performed alternately in the 3DGS training process until convergence, as summarized in Algorithm 1. Generally, we consider our solution convergences once the condition, $\|\bm{a}-\bm{z}\|^{2}\leq\epsilon$, is satisfied, or the iterations reach the predefined maximum number.
Remark-3. In summary, our “optimizing-sparsifying" solution does not explicitly prune the Gaussians but temporarily “marks" them to be pruned on the introduced auxiliary variable $\bm{z}$. As those steps are performed alternately, summarized in Algorithm 1, the process eventually converges to a sparse set that maximizes the performance. Upon the convergence, the information can be preserved and transferred to the “non-zero" Gaussians. Another benefit over existing heuristic one-shot pruning approaches is that our removal of Gaussians does not rely on a one-time importance measurement but a long-term training process. As shown in Figure 2, existing one-shot, criterion-based methods lead to a sudden performance drop, followed by irreversible losses of 0.2 dB and 0.5 dB for importance-based and opacity-based methods, respectively, even with additional training iterations. In contrast, our GaussianSpa approach smoothly and optimally preserves knowledge, enabling more accurate view rendering. Moreover, the introduced variables, $\bm{z}$ and $\bm{\lambda}$, are considered constants, as we do not calculate their gradients. The “sparsifying" step is performed every tens iterations instead of every single iteration. Thus, the overall training speed is very similar to the original 3DGS.
Algorithm 1 Procedure of “Optimizing-Sparsifying"
Output: Optimized $\bm{a}$ and $\bm{\Theta}$.
[…]
2:$t$leftarrow00;
3:while $\|\bm{a}-\bm{z}\|^{2}>\epsilon$ and $t\leq T$ do
4:     Update $\bm{a}$ and $\bm{\Theta}$ via Eq. 14; $\triangleright$ “Optimizing" Step
5:     Update $\bm{z}$ via Eq. 16; $\triangleright$ “Sparsifying" Step
6:     Update $\bm{\lambda}$ via Eq. 17; $\triangleright$ Multiplier Update
[…]
8:end while

### v3

# GaussianSpa: An “Optimizing-Sparsifying" Simplification Framework for Compact and High-Quality 3D Gaussian Splatting

@@ Abstract
3D Gaussian Splatting (3DGS) has emerged as a mainstream for novel view synthesis, leveraging continuous aggregations of Gaussian functions to model scene geometry. However, 3DGS suffers from substantial memory requirements to store the large amount of Gaussians, hindering its efficiency and practicality. To address this challenge, we introduce GaussianSpa, an optimization-based simplification framework for compact and high-quality 3DGS. Specifically, we formulate the simplification objective as a constrained optimization problem associated with the 3DGS training. Correspondingly, we propose an efficient “optimizing-sparsifying” solution for the formulated problem, alternately solving two independent sub-problems and gradually imposing substantial sparsity onto the Gaussians in the 3DGS training process. We conduct quantitative and qualitative evaluations on various datasets, demonstrating the superiority of GaussianSpa over existing state-of-the-art approaches. Notably, GaussianSpa achieves an average PSNR improvement of 0.9 dB on the real-world Deep Blending dataset with 10$\times$ fewer Gaussians compared to the vanilla 3DGS. Our project page is available at https://noodle-lab.github.io/gaussianspa.
Figure 1: We present GaussianSpa, enabling high-quality and compact view synthesis with superior rendering of details. Compared to the existing state-of-the-art method, Mini-Splatting [17], our GaussianSpa captures detail-rich textures and low-frequency background more accurately with fewer Gaussians.
@@ 1 Introduction
Novel view synthesis has become a pivotal area in computer vision and graphics, advancing applications such as virtual reality, augmented reality, and immersive media experiences [18]. NeRF [32] has recently gained prominence in this domain because it can generate high-quality, photorealistic images from sparse input views by representing scenes as continuous volumetric functions based on neural networks. However, NeRF requires substantial computational resources and long training times, making it less practical for real-time and large-scale applications.
@@
3D Gaussian Splatting (3DGS) [22] has emerged as a powerful alternative, leveraging continuous aggregations of Gaussian functions to model scene geometry and appearance. Unlike NeRF, which relies on neural networks to approximate volumetric radiance fields, 3DGS directly represents scenes using a collection of Gaussians. This approach effectively captures details and smooth transitions, offering faster training and rendering. 3DGS achieves superior visual fidelity [25] compared to NeRF while reducing computational overhead, making it more suitable for interactive applications that demand quality and performance.
@@
Despite its strengths, 3DGS suffers from significant memory requirements that hinder its practicality. The main issue is the massive memory consumption required to store numerous Gaussians, each with parameters like position, covariance, and color. In densely sampled scenes, the sheer volume of Gaussians leads to memory usage that exceeds the capacity of typical hardware, making it challenging to handle higher-resolution scenes and limiting its applicability in resource-constrained environments.
@@
Existing works, e.g., Mini-Splatting [17], LightGaussian [16], LP-3DGS [51], EfficientGS [28], and RadSplat [37], have predominantly focused on mitigating this issue by removing a certain number of Gaussians. Techniques such as pruning and sampling aim to discard unimportant Gaussians based on hand-crafted criteria such as opacity [22, 50], importance score (hit count) [17, 16, 37], dominant primitives [28], and binary mask [51, 26]. However, such single-perspective heuristic criteria may lack robustness in dynamic scenes or varying lighting. Moreover, the sudden one-shot removal may cause permanent loss of Gaussians crucial to visual synthesis, making it challenging to recover the original performance after even long-term training, as shown in Figure 2. Mask-based pruning methods also suffer from this issue due to weak sparsity enforcement on the multitude of Gaussians. Consequently, while these methods can alleviate memory and storage burdens to some extent, they often lead to sub-optimal rendering outcomes with loss of details and visual artifacts, thereby compromising the quality of the synthesized views.
Figure 2: PSNR curves of hand-crafted criteria-based pruning methods. Gaussians are removed by 85% at iteration 25K.
@@
In this paper, we present an optimization-based simplification framework, GaussianSpa, for compact and high-quality 3DGS. In the proposed framework, we formulate 3DGS simplification as a constrained optimization problem under a target number of Gaussians. Then, we propose an efficient “optimizing-sparsifying" solution for the formulated problem by splitting it into two simple sub-problems that are alternately solved in the “optimizing" step and the “sparsifying" step. Instead of permanently removing a certain number of Gaussians, GaussianSpa incorporates the “optimizing-sparsifying” algorithm into the training process, gradually imposing a substantial sparse property onto the trained Gaussians. Hence, our GaussianSpa can simultaneously enjoy maximum information preservation from the original Gaussians and a desired number of reduced Gaussians, providing compact 3DGS models with high-quality rendering. Overall, our contributions can be summarized as follows:
@@
  • We comprehensively evaluate GaussianSpa via extensive experiments on various complex scenes, demonstrating improved rendering quality compared to existing state-of-the-art approaches. Particularly, with as high as 10$\times$ fewer number of Gaussians than the vanilla 3DGS, GaussianSpa achieves an average 0.4 dB PSNR improvement on the Mip-NeRF 360 [2] and Tanks&Temples [24] datasets, 0.9 dB on Deep Blending [20] dataset. Furthermore, we conduct various visual quality analyses, demonstrating the superiority of GaussianSpa in rendering high-quality views and capturing rich details.
@@ 2 Related Work
Figure 3: Overall workflow of our proposed GaussianSpa framework.
@@ 2.1 Novel View Synthesis
Novel View Synthesis (NVS) [1] generates images of a 3D scene from unseen viewpoints based on existing multi-view data. A seminal work in this field is NeRF [32], which uses an implicit neural network to represent a scene as a continuous 5-dimensional function, mapping 3D coordinates and viewing directions to color and density. However, its reliance on single narrow-ray sampling poses challenges in addressing aliasing and blur artifacts. More recently, 3D Gaussian Splatting (3DGS), which leverages continuous aggregations of Gaussian functions to model scene geometry, has emerged as a point-based alternative, demonstrating significant improvements in rendering quality and speed.
@@ 2.2 Efficient 3DGS
3DGS significantly improves rendering quality and computational efficiency compared to traditional methods such as NeRF [32], which are often resource-intensive and computationally slow. However, the demands of real-time rendering and the reconstruction of unbounded, large-scale 3D scenes underscore the requirement for model efficiency [7]. Optimization for efficient 3DGS focuses on four main categories: Gaussian simplification, quantization, spherical harmonics optimization, and hybrid compression.
@@
Gaussian Simplification. Gaussian simplification involves removing Gaussians to compress 3DGS models into compact forms, significantly reducing memory costs. Prior approaches [17, 37, 31, 51, 29, 43, 27, 10, 23] mainly focus on pruning redundant Gaussians according to hand-crafted importance criteria. Mini-Splatting [17] addresses overlapping and reconstruction issues with blur splitting, depth reinitialization, and stochastic sampling. Radsplatting [37] enhances robustness by using a max operator to compute importance scores from ray contributions. Taming 3DGS [31] applies selective densification based on pixel saliency and gradients, while LP-3DGS [51] learns a binary mask for efficient pruning.
@@
Quantization. Quantization [47] aims to compress data using discrete entries without appreciable degradation in quality, reducing data redundancy. Several works [19, 35, 9, 48, 39, 33, 36] apply quantization for compressing 3DGS models. For example, EAGLES [19] quantizes attributes like spherical harmonics (SH) coefficients. Vector quantization, a form of quantization, compresses data by mapping vectors to a learned codebook. CompGS [35] leverages K-means-based vector quantization to reduce storage and rendering time, while RDO-Gaussian [48] combines pruning and entropy-constrained vector quantization for efficient compression with minimal quality loss.
@@
Spherical Harmonics Optimization. Among 3DGS parameters, spherical harmonics (SH) coefficients take up a significant portion, requiring (45+3) floating-point values per Gaussian for a degree of 3 [22], which accounts for around 80% of the total attribute volume. Prior works [16, 26] optimize SH coefficients to reduce storage overhead. For instance, EfficientGS [28] retains zeroth-order SH for all Gaussians, increasing the order when necessary.
@@
Hybrid Compression. In addition to methods focused on individual aspects of 3DGS compression, several recent approaches [16, 28, 26, 49] combine multiple techniques for hybrid compression to achieve further storage reduction. LightGaussian [16] uses a global importance score for pruning and applies vector quantization to compress SH coefficients. EfficientGS [28] selectively increases SH order and densifies non-steady Gaussians. CompactGaussian [26] introduces a learnable mask for pruning, uses residual vector quantization for scale and rotation, and replaces SH with a grid-based neural field for color representation.
@@
Although mask-based method impose sparsity relying on mask loss, they and other methods both lead to irreversible performance loss in the Gaussian simplification process by performing a one-shot pruning on the Gaussians, as shown in Figure 2. Consequently, performance often worsens when applying SH optimization and quantization due to accumulated errors. Our work focuses on Gaussian simplification by formulating it as an optimization problem and imposing substantial sparsity onto Gaussians, mitigating performance issues with improved rendering quality.
@@ 3.1 Background of 3D Gaussian Splatting
where $\bm{x}$ is an arbitrary location in the 3D scene. The covariance matrix $\bm{\Sigma}$ is generally composed of $\bm{\Sigma}=\bm{R}\bm{S}\bm{S}^{\top}\bm{R}^{\top}$ with a rotation matrix $\bm{R}$ and a scale matrix $\bm{S}$, ensuring the positive definite property. Each Gaussian also has an associated opacity $a$ and spherical harmonics (SH) coefficients.
@@
In the training process [22], the Gaussians are initialized from a sparse point cloud generated by Structure-from-Motion (SfM) [44]. Then, the number of Gaussians is adjusted by a densification algorithm [22] according to the reconstruction geometry. Afterward, each Gaussian’s attributes, including the center position, rotation matrix, scaling matrix, opacity, and SH coefficients, are optimized by minimizing the reconstruction errors using the standard gradient descent. Specifically, the loss $\mathcal{L}$ is given by
$\mathcal{L}=(1-\rho)\mathcal{L}_{1}+\rho\mathcal{L}_{\text{D-SSIM}},\vspace{-1mm}$ (4)
@@ 3.2 Problem Formulation: Simplification as Constrained Optimization
However, the above problem cannot be solved because the restriction on the number of Gaussian functions is not differentiable. Recalling the rendering process with Eq. 3, the contribution of the $i$-th Gaussian to a particular pixel is finally determined by its opacity $a_{i}$. Alternatively, the original sparsity constraint directly on the Gaussians can be explicitly transformed into a sparsity constraint on the opacity variables. As each Gaussian only has one opacity variable, we can use a vector $\bm{a}$ to represent all the opacities and easily restrict its number of non-zeros with $\ell_{0}$ norm, $\|\bm{a}\|_{0}$. For better presentation, we use a variable set $\bm{\Theta}$ to represent other GS variables except for the opacity $\bm{a}$. Hence, the constrained optimization problem can be reformulated as
@@
Algorithm 1 Procedure of “Optimizing-Sparsifying"
Output: Optimized $\bm{a}$ and $\bm{\Theta}$.
[…]
2:$t$leftarrow00;
3:while $\|\bm{a}-\bm{z}\|^{2}>\epsilon$ and $t\leq T$ do
4:     Update $\bm{a}$ and $\bm{\Theta}$ with Eq. 14; $\triangleright$ “Optimizing" Step
5:     Update $\bm{z}$ with Eq. 16; $\triangleright$ “Sparsifying" Step
6:     Update $\bm{\lambda}$ with Eq. 17; $\triangleright$ Multiplier Update
[…]
8:end while
@@
$h(\bm{a})=\left\{\begin{matrix}0,&\|\bm{a}\|_{0}\leq\kappa,\\ +\infty,&\textup{otherwise},\\ \end{matrix}\right.\vspace{-1mm}$ (7)
@@
$\min_{\bm{a},\bm{\Theta}}~{}\mathcal{L}(\bm{a},\bm{\Theta})+h(\bm{a}).\vspace{-1mm}$ (8)
Here, the first term is the original 3DGS training objective, which is easy to handle, while the second term is still non-differentiable. This prevents the above problem from being solved using gradient-based solutions under existing training frameworks such as PyTorch [41]. To decouple the two terms, we further introduce an auxiliary variable $\bm{z}$, whose size is identical to $\bm{a}$, and convert Eq. 8 as
@@
respectively. Here, $\eta$ is the learning rate during training.
@@
[…]
$\bm{z}\leftarrow\mathrm{\bf{prox}}_{h}(\bm{a}+\bm{\lambda}).$ (16)
The proximal operator maps the opacity $\bm{a}$ to the auxiliary variable $\bm{z}$, which contains at most $\kappa$ non-zeros. This operator results in a projection solution [5], sparsifying $(\bm{a}+\bm{\lambda})$, by projecting it onto a set that satisfies $\|\bm{z}\|_{0}\leq\kappa$. More specifically, the projection keeps top-$\kappa$ elements and sets the rest to zeros.
[…]
$\bm{\lambda}\leftarrow\bm{\lambda}+\bm{a}-\bm{z}.$ (17)
Overall, the “optimizing," “sparsifying," and multiplier update steps are performed alternately in the 3DGS training process until convergence, as summarized in Algorithm 1. Generally, we consider our solution convergences once the condition, $\|\bm{a}-\bm{z}\|^{2}\leq\epsilon$, is satisfied, or the iterations reach the predefined maximum number.
Remark. The “optimizing" step optimizes over the original Gaussian variables and pushes the Gaussian opacity $\bm{a}$ close to the exactly sparse auxiliary variable $\bm{z}$ using a gradient-based approach, simultaneously satisfying the 3DGS performance and the sparsity requirements. The “sparsifying" step essentially prunes the auxiliary variable $\bm{z}$, which can be considered the exactly sparse version of $\bm{a}$. As those steps are performed alternately, summarized in Algorithm 1, the process eventually converges to a sparse set that maximizes the performance. Upon the convergence, most Gaussians become “zeros," and the information can be preserved and transferred to the “non-zero" Gaussians.